A random number generator gives results that are uniformly distributed with a minimum of 43 and a maximum of 70. What is the probability that a random number is between 45 and 58?
Answer #1

Solution :

Given that,

a = 43

b = 70

P(c < x < d) = (d - c) / (b - a)

P(45< x < 58) = (58 - 45) / (70 - 43)=0.4815

probability=0.4815